If   you have a 28 g sample of N2 gas in a 0.5 L flask at a temperature of 29 C.  What must be the pressure.


and



A gas occupies a volume of 584 mL at a pressure of 770 mL of Hg.  When the pressure is changed, the volume becomes 603 mL  If the temperature stays constant, what is the new pressure in mm of Hg.




1
Expert's answer
2021-06-16T05:32:26-0400

1. We use for the first question Mendeleev Klapiron's formula

pV=(m/M)RT

R=8.31J/mol*K

T=273+29=302 K

V=0.5L

p=mRT/MV=28g*8.31*302K/28g/mol*0.5L=5019.24=5.02 kPa

2. New pressure is 795.05 ml of Hg.